Suppose there are 1000 students in a business administration program. 620 students are in an IT class, 450 students are in a Statistics class, 200 students are taking both.

Create a contingency table for the data above.
Answer #1
IT Non-IT Total
Statistics 200 450-200=250 450
Non-Statistics 420 130 1000-450=550
Total 620 1000-620=380 1000

Bold letters are given.